Which equation below represents a line that is perpendicular to the equation 4x – 12y = 2 and passes through the point (10, –1)?

Answer by mananth(16946)

4 x -12 y = 2
Find the slope of this line

-12 y = -4 x + 2
Divide by -12
y = 1/ 3 x + - 1/6
Compare this equation with y=mx+b
slope m = 1/3

The slope of a line perpendicular to the above line will be the negative reciprocal -3
m1*m2=-1


m= -3 ,point ( 10 , -1 )
Find b by plugging the values of m & the point in
y=mx+b
-1 = -30.00 + b
b= 29
m= -3
The required equation is y = -3 x + 29
m.ananth@hotmail.ca